OpenStudy (anonymous):

a building that is 250 feet cast a shadow 40 feet long. Find the angle of elevation to the nearest tenth of a degree of the sun at this time?

OpenStudy (anonymous):



OpenStudy (anonymous):

apply tangent to get the value of theta....

OpenStudy (anonymous):

how would i set up the equation

OpenStudy (anonymous):

well u should know the trigonometry ratios.... since \[\tan \theta = \frac{ perpendicula }{ base }\] in present case we have perpendicular 250ft & base = 40ft put these values..you will get ur answer

OpenStudy (anonymous):

okay well i dont use perpendicular and base i go by opposite and adjacent!

OpenStudy (anonymous):

well in that case \[\tan \theta = \frac{ opposite }{ adjacent }\] and since adjacent is next to theta which is 40ft while opposite is opposite to the angle theta which is 250ft.. put these values and solve u will get ur answer

OpenStudy (anonymous):

\[\tan 40/250\]

OpenStudy (anonymous):

is that what i should plug into the calculator

OpenStudy (anonymous):

nope ...since adjacent is that side which is next to theta so it means in present case adjacent = 40ft while opposite is 250ft so equation should be \[\tan \theta = \frac{ 250 }{ 40 }\]

OpenStudy (anonymous):

i got 0.109.... which i don't think is correct

OpenStudy (anonymous):

and for finding the value of theta u should take the inverse... then equation will become \[\theta = \tan^{-1} \frac{ 250 }{ 40 }\]

OpenStudy (anonymous):

ur answer is wrong because ur taking the tan of (250/40)...while u should take the tan inverse of (250/40)...

OpenStudy (anonymous):

like i have mentioned above...

OpenStudy (anonymous):

i did it again and got 80.9097.....

OpenStudy (anonymous):

yup...this is the right answer..

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

n.p